How many gallons of a 70% acid solution must be mixed with 70 gallons of a 30% solution to obtain a solution that is 50% acid?
Answer by jorel1380(3719)

.7n+70(.30)=.5(n+70)
.7n+21=.5n+35
.2n=14
n=14/.2=70
You need 70 gallons of 70% solution to make the required mix..
70(.7)+70(.3)=70
140(.5)=70